You invested $7000 in two accounts paying 7% and 6% annual interest, respectively. If the total interest earned for the year was $433, Let x denote the amount of money invested at 7%. Find an expression in terms of x for the amount invested at 6%.
Answer by richwmiller(17219)

x+y=7000,
0.07*x+0.06*y=433
x=7000-y
0.07*(7000-y)+0.06*y=433
490-0.07y+0.06*y=433
-0.01*y=-57
y= 5700
x=7000-y
x=1300
check
0.07*1300+0.06*5700=433
91+342=433
433=433
ok